OpenStudy (anonymous):

Given that Sinθ=5Cosθ, (∴ Tanθ=5) Hence find the values of θ in the interval 0≤x≤360 for which Sinθ=5Cosθ

OpenStudy (eyust707):



OpenStudy (anonymous):

in otherwords, find values of θ for which Tanθ=5 \[\theta=tan^{-1}(5)\] \[\theta=78.69^o\] trig functions are periodic, and tanθ=5 each: \[78.69+n\pi\] where n is integer.

OpenStudy (anonymous):

n only equal to 0 and 1

OpenStudy (anonymous):

nope :P, not only 0 and 1, and integer, -1,-2,-10000000000, or 10000000 answer will be 5

OpenStudy (anonymous):

any*

OpenStudy (anonymous):

0<x<360

OpenStudy (anonymous):

oh, oops

OpenStudy (anonymous):

if your n=100, then it will be 100pi+78.69=3600+78.69